# The Kobayashi-Hitchin correspondence for nef and big classes

### 11 Mar 2026 v1

@@ 3.2.1 Reduction of the construction to a uniform $C^{0}$-estimate
holds for any $i$, then there exists a holomorphic morphism $\Psi_{\infty}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}:=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ such that the following holds:
@@ 1. $(1)$
    there exists a subsequence of $\{\Psi_{i}\}_{i}$, also denoted by $\{\Psi_{i}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$,
    $\lim_{i\to\infty}\int_{A}|\Psi_{i}-\Psi_{\infty}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
@@ 2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that
    $\sup_{X}|s_{D}\Psi_{\infty}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
@@ 3. $(3)$
Let $\eta_{k}:X\to[0,1]$ be a sequence of smooth functions such that ${\rm{Supp}}(\eta_{k})\Subset X\setminus Z$ and
$A_{k}:=\{\eta_{k}=1\}\nearrow X\setminus Z$ (3.2)
in $k\to\infty$. Such $\eta_{k}$ exists by Lemma 3.2.
Let us define an $h_{0}^{*}\otimes h_{0}\otimes h_{D}$-connection on ${\rm{End}}(E)\otimes\mathcal{O}(D)$ by $\widetilde{\nabla_{i}}:=\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}$. Since $\Psi_{i}:(E,\bar{\partial}^{E})\to(E,\bar{\partial}^{\nabla_{i}})$ is holomorphic, we immediately have $\bar{\partial}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})=0$. Then if we use the identity $\Box^{\partial^{\widetilde{\nabla_{i}}}}_{\omega_{i}}=\Box^{\bar{\partial}^{\widetilde{\nabla_{i}}}}_{\omega_{i}}-\sqrt{-1}[\Lambda_{\omega_{i}},F_{\widetilde{\nabla_{i}}}$], we obtain
$\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})$ $=2\Box_{\omega_{i}}^{\bar{\partial}^{\widetilde{\nabla_{i}}}}(s_{D}\Psi_{i})-\sqrt{-1}\Lambda_{\omega_{i}}F_{\widetilde{\nabla_{i}}}\cdot(s_{D}\Psi_{i})$
$=-\sqrt{-1}\Lambda_{\omega_{i}}F_{\widetilde{\nabla_{i}}}\cdot(s_{D}\Psi_{i}).$
In particular, we obtain
$\|\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})\|_{L^{2}(A_{k})}\leq C_{k}$ (3.3)
for any $i$. Next we show
$\|\widetilde{\nabla_{i}}(\eta_{k}s_{D}\Psi_{i})\|^{2}_{L^{2}(X)}\leq(C^{\prime}_{k}+1)\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}+\|\Delta_{\omega_{i}}^{\widetilde{\nabla_{i}}}(s_{D}\Psi_{i})\|_{L^{2}(A_{2k})}^{2}\leq C_{k}.$ (3.4)
First we compute as follows:
$(\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i}),\eta_{k}^{2}s_{D}\Psi_{i})_{L^{2}}$ $=(\widetilde{\nabla_{i}}(\eta_{k}^{2}s_{D}\Psi_{i}),\widetilde{\nabla_{i}}s_{D}\Psi_{i})$
$=(2\eta_{k}\nabla\eta_{k}\cdot s_{D}\Psi_{i}+\eta_{k}^{2}\widetilde{\nabla_{i}}(s_{D}\Psi_{i}),\widetilde{\nabla_{i}}(s_{D}\Psi_{i}))$
$=2(s_{D}\Psi_{i}\nabla\eta_{k},\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i}))+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$\geq-2(2\|s_{D}\Psi_{i}\nabla\eta_{k}\|)(\frac{1}{2}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|)+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$\geq-2(4\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}+\frac{1}{4}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2})+\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$
$=-8\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}+\frac{1}{2}\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}.$
Thus, we obtain
$\|\eta_{k}\widetilde{\nabla_{i}}(s_{D}\Psi_{i})\|^{2}$ $\leq 2(\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i}),\eta_{k}^{2}s_{D}\Psi_{i})+16\|s_{D}\Psi_{i}\nabla\eta_{k}\|^{2}$
$\leq 2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|_{L^{2}(A_{2k})}\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}+C_{k}\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}$
$\leq(1+C_{k})\|s_{D}\Psi_{i}\|_{L^{2}(A_{2k})}^{2}+\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}.$
Therefore, by the assumption of this lemma and with (3.3), the following estimate follows:
$\|\widetilde{\nabla_{i}}(\eta_{k}s_{D}\Psi_{i})\|^{2}$ $\leq 2(1+C^{\prime}_{k})\|s_{D}\Psi_{i}\|^{2}_{L^{2}(A_{2k})}+2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}\leq C^{\prime\prime}_{k},$
which shows (3.4). By Lemma 3.5 (5), $\nabla_{i}$ smoothly converges to $\nabla_{\infty}$ on each $A_{k}$ and by the assumption that $\omega_{i}$ smoothly converges to $T$ on each $A_{k}$. Therefore, if we define $\widetilde{\nabla_{\infty}}:=\nabla_{\infty}\otimes\nabla_{h_{0}}\otimes\nabla_{h_{D}}$, then we obtain that
$\|\widetilde{\nabla_{\infty}}(\eta_{k}s_{D}\Psi_{i})\|^{2}_{L^{2}(X)}\leq C_{k}\|s_{D}\Psi_{i}\|^{2}_{L^{2}(A_{2k})}+2\|\Delta^{\widetilde{\nabla_{i}}}_{\omega_{i}}(s_{D}\Psi_{i})\|^{2}_{L^{2}(A_{2k})}\leq C^{\prime}_{k},$ (3.5)
By the assumption (3.1) and above (3.5), the Sobolev embedding theorem ensures that, for any $k$, we can find a subsequence $\{i^{k}_{j}\}_{j}$ of $\{i\}_{i}$ such that $\{i^{k+1}_{j}\}_{j}\subset\{i^{k}_{j}\}_{j}$ and $\eta_{k}\Psi_{i^{k}_{j}}\xrightarrow{j\to\infty}\widetilde{\Psi_{k}}$ strongly in $L^{2}(X)$. Since $\eta_{k}=\eta_{k+1}=\cdots$ on $A_{k}=\{\eta_{k}=1\}$ by (3.2) and by the choice of the subsequence, we can assume that $\widetilde{\Psi_{k}}=\widetilde{\Psi_{k+1}}$$\cdots$ on $A_{k}$. Then we define $\Psi_{\infty}$ by $\Psi_{\infty}(x):=\widetilde{\Psi_{k}}(x)$ for $x\in A_{k}$. Then $\Psi_{\infty}$ is well-defined and it is contained in $L^{2}_{{\rm{loc}}}(X\setminus Z)$. By the diagonal argument, we can find a subsequence $\{i^{k}_{j_{k}}\}_{k}$ of $\{i\}_{i}$ such that $\Psi_{i^{k}_{j_{k}}}\xrightarrow{k\to\infty}\Psi_{\infty}$ in $L^{2}_{{\rm{loc}}}(X\setminus Z)$. Since $\Psi_{i^{k}_{j_{k}}}$ converges to $\Psi_{\infty}$ at almost everywhere points in $X\setminus Z$ , the assumption (3.1) ensures that
$\sup_{X}|s_{D}\Psi_{\infty}|_{h_{0},h_{D}}\leq C.$
Since $\bar{\partial}^{{\nabla_{i}}\otimes\nabla_{h_{0}}^{*}}\Psi_{i}=0$, we can easily see that $\bar{\partial}^{\nabla_{\infty}\otimes\nabla_{h_{0}}^{*}}\Psi_{\infty}=0$ on $X\setminus Z$ in the sense of currents. Since $\Psi_{\infty}$ is an $L^{2}$ local section on $X\setminus Z$, the standard convolution argument shows that $\Psi_{\infty}$ is holomorphic with respect to $\bar{\partial}^{{\nabla_{\infty}}\otimes\nabla_{h_{0}}^{*}}$ (e.g. [Jin25-2, proof of Proposition 4.15]). Finally we see that $\Psi_{\infty}$ is nontrivial. Since $\det\Psi_{i^{k}_{j_{k}}}\equiv 1$, we have $|\Psi_{i^{k}_{j_{k}}}|\geq 1$. Since $\Psi_{i^{k}_{j_{k}}}$ converges to $\Psi_{\infty}$ almost everywhere on $X\setminus Z$, we have $|\Psi_{\infty}|\geq 1$ almost everywhere on $X\setminus Z$. Since $\Psi_{\infty}$ is continuous on $X\setminus Z$, we have that $|\Psi_{\infty}|\geq 1$ on $X\setminus Z$. In particular, we have that $\Psi_{\infty}$ is nontrivial. ∎
@@ 3.2.4 Construction of $\Psi_{\infty}$
As a consequence of the previous subsections, we obtain the followings
@@
We use the notations Notation 3.2 and Lemma 3.5. Then there exists a holomorphic morphism ${\Psi_{\infty}}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ which satisfies the following conditions:
@@ 1. $(1)$
    There exists a subsequence of $\{{\Psi_{i}}\}_{i}$, also denoted by $\{{\Psi_{i}}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$, the following holds:
    $\lim_{i\to\infty}\int_{A}|{\Psi_{i}}-{\Psi_{\infty}}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
@@ 2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that the following holds for any $i$:
    $\sup_{X}|s_{D}{\Psi_{i}}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
@@ 3. $(3)$
By Lemma 3.8 and Lemma 3.9, we can see that the same argument with Theorem 3.13 applies to $\Psi_{i}^{-1}$, and thus we obtain the following theorem:
@@
We use the notations Notation 3.2 and Lemma 3.5. Then there exists a holomorphic morphism $\Phi_{\infty}:E_{\infty}=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ which satisfies the following conditions:
@@ 1. $(1)$
    There exists a subsequence of $\{\Psi_{i}^{-1}\}_{i}$, also denoted by $\{\Psi_{i}^{-1}\}_{i}$, such that for any compact subset $A\subset X\setminus Z$, the following holds:
    $\lim_{i\to\infty}\int_{A}|\Psi_{i}^{-1}-\Phi_{\infty}|^{2}_{h_{0}}\omega_{i}^{n}=0.$
@@ 2. $(2)$
    There exists a constant $C>0$ independent of $i$ such that the following holds for any $i$:
    $\sup_{X}|s_{D}\Psi_{i}^{-1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
  3. $(3)$
    Furthermore, $\Phi_{\infty}$ is nontrivial: $\Phi_{\infty}\not\equiv 0$.
As a corollary of Theorem 3.13 and Theorem 3.14, we can derive the following uniform estimate on the first derivatives.
Proposition 3.15.
We use notations in Notation 3.2. Let $\Psi_{i}$}, $\Psi_{\infty}$}, $\Phi_{\infty}$ be holomorphic morphisms constructed in Theorem 3.13 and Theorem 3.14. There exists a constant $C>0$ such that the following inequality holds for any $i$:
$\int_{X}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}{\Psi_{i}}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},\omega_{i}}\omega_{i}^{n}\leq C.$
The same estimate holds for $\Psi_{i}^{-1}$. In particular, we have
$\int_{X\setminus Z}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}{\Psi_{\infty}}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},T}T^{n}<\infty.$
The same estimate holds for $\Phi_{\infty}$.
@@ 3. $(3)$
Let $h_{D}$ be a smooth hermitian metric on $\mathcal{O}(D)$. We denote by $\nabla_{D}:=\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}$ a $h_{0}^{*}\otimes h_{0}\otimes h_{D}$-connection on ${\rm{Hom}}(E,E)\otimes\mathcal{O}(D)$. If we use the identity $\Box^{\partial^{\nabla_{D}}}_{\omega_{i}}=\Box^{\bar{\partial}^{\nabla_{D}}}_{\omega_{i}}-\sqrt{-1}[\Lambda_{\omega_{i}},F_{\nabla_{D}}]$, and $T\geq|s_{D}|^{2m}\omega_{0}$, the following calculation works:
$\leq C.$ (3.34)
Since $\sup_{X\setminus D}|s_{D}\Psi_{i}|\leq C$ by Theorem 3.13, we obtain that
$=\Psi_{i}\circ(\nabla_{h_{i}}-\nabla_{h_{0}})$
$=\Psi_{i}\circ(\Psi_{i}^{-2}\partial^{\nabla_{h_{0}}}(\Psi_{i}^{2}))$
$=2\partial^{\nabla_{h_{0}}}\Psi_{i}.$
3.3 Existence of $T$-adapted HYM metrics
Although this subsection overlaps with [Jin25-2, §4.2], we include the proof for the completeness.
Notation 3.16.
We use the notations in Notation 3.2. Let $Z=\overline{D\cup\Sigma}\subset X$ be the compact subset of Hausdorff codimension 2 and $\nabla_{\infty}$ be the $T$-admissible HYM connection of the complex hermitian vector bundle $(E,h_{0})|_{X\setminus Z}$ in Lemma 3.5. Let ${\Psi_{\infty}}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}:=(E|_{X\setminus Z},\bar{\partial}^{\nabla_{\infty}})$ and $\Phi_{\infty}:E_{\infty}\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ be holomorphic morphisms constructed in Theorem 3.13 and Theorem 3.14.
Theorem 3.17.
We use notations in Notation 3.16. Then the following holds.
  1. $(1)$
    $\Psi_{\infty}:(E,\bar{\partial}^{E})|_{X\setminus Z}\to E_{\infty}$ and $\Phi_{\infty}:E_{\infty}\to(E,\bar{\partial}^{E})|_{X\setminus Z}$ satisfies $\Psi_{\infty}\circ\Phi_{\infty}={\rm{id}}$ and $\Phi_{\infty}\circ\Psi_{\infty}={\rm{id}}$. That is, $\Phi_{\infty}=\Psi_{\infty}^{-1}$. In particular $\Psi_{\infty}$ is a holomorphic isomorphism.
  2. $(2)$
    A smooth hermitian metric $h_{\infty}:=h_{0}{\Psi_{\infty}}^{2}$ on $E|_{X\setminus Z}$ is a $T$-adapted HYM metric on $E$ whose Chern connection is ${\Psi_{\infty}}^{-1}\circ\nabla_{\infty}\circ{\Psi_{\infty}}$.
Proof.
(1) By Theorem 3.13, we have $\lim_{i\to\infty}\Psi_{i}=\Psi_{\infty}$ almost everywhere on $X\setminus Z$, and by Theorem 3.14, we have $\lim_{i\to\infty}\Psi_{i}^{-1}=\Phi_{\infty}$ almost everywhere on $X\setminus Z$. Since $\Psi_{\infty}$ and $\Phi_{\infty}$ are both holomorphic, we obtain $\Psi_{\infty}\circ\Phi_{\infty}={\rm{id}}$ and $\Phi_{\infty}\circ\Psi_{\infty}={\rm{id}}$ on $X\setminus Z$ and thus $\Psi_{\infty}$ is a holomorphic isomorphism.
(2) By direct computation, we can see that ${\Psi_{\infty}}^{-1}\circ\nabla_{\infty}\circ{\Psi_{\infty}}$ is the Chern connection of $h_{\infty}$. By Theorem 3.13, it suffices to prove that the curvature tensor $F_{h_{\infty}}$ is $L^{2}$-integrable with respect to the norm defined by $h_{\infty}$ and $T$. We remark that $F_{h_{\infty}}={\Psi_{\infty}}^{-1}\circ F_{\nabla_{\infty}}\circ{\Psi_{\infty}}$ and $(\sqrt{-1}F_{h_{\infty}})^{*_{h_{\infty}}}=\sqrt{-1}F_{h_{\infty}}$. Then, the point-wise identity $|F_{h_{\infty}}|_{h_{\infty},T}^{2}T^{n}={\rm{Tr}}(\sqrt{-1}F_{h_{\infty}}\wedge\sqrt{-1}F_{h_{\infty}})\wedge T^{n}+|\Lambda_{T}F_{h_{\infty}}|_{h_{\infty}}^{2}T^{n}$ on $X\setminus Z$ (c.f. [Kob87] Chapter 4 section 4) shows $|F_{h_{\infty}}|_{h_{\infty},T}^{2}=|F_{\nabla_{\infty}}|_{h_{0},T}^{2}$ on $X\setminus Z$. Thus $h_{\infty}$ is a $T$-adapted HYM metric on $E|_{X\setminus Z}$. ∎
In conclusion, we obtain the following theorem.
@@ 2. $(2)$
    $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi|^{2}T^{n}\leq C.$
Lemma 3.19.
We use notations in Notation 3.16. Let $h_{\infty}:=h_{0}\Psi_{\infty}^{2}$ be the $T$-adapted HYM metric on $E$ constructed in Theorem 3.17. Then the $T$-HYM constant $\lambda$ of $h_{\infty}$ satisfies the following:
$\lambda=\frac{1}{\alpha^{n}}\frac{\int_{X\setminus D}\sqrt{-1}{\rm{Tr}}F_{h_{\infty}}\wedge T^{n-1}}{{\rm{rk}}E}=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
Proof.
The first equality in the statement directly follows from the $T$-HYM equation of $h_{\infty}$. We show the second equation. Let $\nabla_{i}$ be the $\omega_{i}$-HYM connection on $(E,h_{0})$ in Lemma 3.5 and $\nabla_{\infty}:=\Psi_{\infty}\circ\nabla_{h_{\infty}}\circ\Psi_{\infty}^{-1}$ be the $T$-admissible HYM connection on $(E,h_{0})$ such that $\nabla_{i}\to\nabla_{\infty}$ in $C^{\infty}_{{\rm{loc}}}(X\setminus Z)$ as in Lemma 3.5. It is obvious that the $T$-HYM constant of $\nabla_{\infty}$ coincides with that of $h_{\infty}$. Since $\sqrt{-1}\Lambda_{\omega_{i}}F_{\nabla_{i}}\to\sqrt{-1}\Lambda_{T}F_{\nabla_{\infty}}$ at each point in $X\setminus D$ and since $\lambda_{i}$, the $\omega_{i}$-HYM constant of $\nabla_{i}$, satisfies $\lambda_{i}=\frac{1}{(\alpha+(1/i)\omega_{0})^{n}}\frac{c_{1}(E)\cdot(\alpha+(1/i)\omega_{0})^{n-1}}{{\rm{rk}}E}$, we obtain

### 18 Mar 2026 v2

@@ 3.2.1 Reduction of the construction to a uniform $C^{0}$-estimate
Lemma 3.7.
We use notations in Notation 3.2. If there exists a constant $C>0$ independent of $i$ such that
$\sup_{X}|s_{D}\Psi_{i}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C$
holds for any $i$. Then there exists a constant $C>0$ such that the following inequality holds for any $i$:
  1. $(1)$
    $\sup_{X}|s_{D}\Psi_{i}^{-1}|_{h_{0}\otimes h_{0}^{*}\otimes h_{D}}\leq C$.
  2. $(2)$
    $\int_{X}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}{\Psi_{i}}^{\pm 1}|^{2}_{h_{0}\otimes h_{0}^{*}\otimes h_{D}^{m+1},\omega_{i}}\omega_{i}^{n}\leq C.$
Proof.
(1) is a direct consequence of $\det\Psi_{i}=1$. We prove (2). Let $h_{D}$ be a smooth hermitian metric on $\mathcal{O}(D)$. We denote by $\nabla_{D}:=\nabla_{i}\otimes\nabla_{h_{0}}^{*}\otimes\nabla_{h_{D}}$ a $h_{0}^{*}\otimes h_{0}\otimes h_{D}$-connection on ${\rm{Hom}}(E,E)\otimes\mathcal{O}(D)$. If we use the identity $\Box^{\partial^{\nabla_{D}}}_{\omega_{i}}=\Box^{\bar{\partial}^{\nabla_{D}}}_{\omega_{i}}-\sqrt{-1}[\Lambda_{\omega_{i}},F_{\nabla_{D}}]$, and $T\geq|s_{D}|^{2m}\omega_{0}$, the following calculation works:
$\leq C.$ (3.1)
Since $\sup_{X\setminus D}|s_{D}\Psi_{i}|\leq C$ by the assumption, we obtain that
$=\Psi_{i}\circ(\nabla_{h_{i}}-\nabla_{h_{0}})$
$=\Psi_{i}\circ(\Psi_{i}^{-2}\partial^{\nabla_{h_{0}}}(\Psi_{i}^{2}))$
$=2\partial^{\nabla_{h_{0}}}\Psi_{i}.$
The statement for $\Psi_{i}^{-1}$ is proved by applying the same argument to the holomorphic morphism $\Psi_{i}^{-1}:(E,\bar{\partial}^{\nabla_{i}})\to(E,\bar{\partial}^{E})$. ∎
@@ 2. $(2)$
holds for any $i$, then there exists a smooth $h_{0}$-hermitian endomorphism $\Psi_{\infty}$ of $E|_{X\setminus D}$ such that the following holds:
@@ 1. $(1)$
    For any $0<\alpha<1$, there exists a subsequence of $\{\Psi_{i}\}_{i}$, also denoted by $\{\Psi_{i}\}_{i}$, such that for any compact subset $A\subset X\setminus D$,
    $\lim_{i\to\infty}\|\Psi_{i}-\Psi_{\infty}\|_{C^{1,\alpha}(A)}=0,$
    where the $C^{1,\alpha}$-norm is defined by the reference smooth Kähler metric $\omega_{0}$.
@@ 2. $(2)$
    There exists a constant $C>0$ such that
    • $\sup_{X\setminus D}|s_{D}\Psi_{\infty}^{\pm 1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
    • $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi_{\infty}^{\pm 1}|^{2}_{h_{D}\otimes h_{0}\otimes h_{0}^{*},T}T^{n}\leq C.$
@@ 3. $(3)$
It is a direct consequence of [BS94, Proposition 1]. In fact, the $\omega_{i}$-HYM equation is locally written as
$\sqrt{-1}\Lambda_{\omega_{i}}\bar{\partial}(\partial h_{i}\cdot h_{i}^{-1})=\lambda_{i}{\rm{id}}.$ (3.3)
By Lemma 3.7, we have
$\sup_{X\setminus D}|s_{D}\Psi_{i}^{-1}|\leq C.$ (3.4)
By (3.2), (3.3) and (3.4), [BS94, Proposition 1] implies that for any relatively compact open subset $K\subset X\setminus D$ and $0<\alpha^{\prime}{}<1$, there exists a constant $C_{K}>0$ such that the following holds:
$\|\Psi_{i}\|_{C^{1,\alpha^{\prime}{}}(K)}\leq C_{K}.$
Here the $C^{1,\alpha^{\prime}{}}$-norm is defined by $\omega_{0}$. Since the natural inclusion $C^{1,\alpha^{\prime}{}}(\overline{K})\hookrightarrow C^{1,\alpha}(\overline{K})$ is compact for any $0<\alpha<\alpha^{\prime}{}<1$, we can find a subsequence of $\{\Psi_{i}\}_{i}$ such that $\Psi_{i}\to\Psi_{\infty,K}$ in $C^{1,\alpha}(K)$. Then, if we consider a sequence of relatively compact open subsets by $X\setminus D_{\varepsilon_{k}}$ for $\varepsilon_{k}\searrow 0$ where $D_{\varepsilon_{k}}$ is the $\varepsilon_{k}$-neighborhood of $D$ with respect to $\omega_{0}$, then the standard diagonal procedure gives a subsequence of $\{\Psi_{i}\}_{i}$ such that $\Psi_{i}\to\Psi_{\infty}$ in $C^{1,\alpha}(K)$ for any relatively compact open subset $K\subset X\setminus D$. Let us define a hermitian metric on $E|_{X\setminus D}$ by $h_{\infty}:=h_{0}\Psi_{\infty}^{2}$. Then it satisfies
$\sqrt{-1}\Lambda_{T}\bar{\partial}(\partial h_{\infty}\cdot h_{\infty}^{-1})=\lambda{\rm{id}}$ (3.5)
where $\lambda=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}$. Since $h_{\infty}^{\pm 1}$ is in $C^{1,\alpha}(K)\cup L^{2}_{1}(K)$ for any $K\Subset X\setminus D$, the standard elliptic estimate concludes that $\Psi_{\infty}$ is smooth on $X\setminus D$. The remaining statements $(2)$ and $(3)$ are the direct consequences of Lemma 3.7, (3.2) and $\det\Psi_{i}=1$. ∎
@@ 3.2.4 Construction of $\Psi_{\infty}$
As a consequence of the previous subsections (refer to Lemma 3.7, Lemma 3.8, Lemma 3.9 and Lemma 3.10 ), we obtain the followings:
@@
We use the notations in Notation 3.2 and Lemma 3.5. Then there exists a smooth $h_{0}$-hermitian endomorphism $\Psi_{\infty}$} of $E$ which satisfies the following conditions:
@@ 1. $(1)$
    For any $0<\alpha<1$, there exists a subsequence of $\{{\Psi_{i}}\}_{i}$, also denoted by $\{{\Psi_{i}}\}_{i}$, such that for any compact subset $A\subset X\setminus D$, the following holds:
    $\lim_{i\to\infty}\|{\Psi_{i}}-{\Psi_{\infty}}\|_{C^{1,\alpha}(A)}=0.$
@@ 2. $(2)$
    There exists a constant $C>0$ such that the following holds:
    • $\sup_{X\setminus D}|s_{D}{\Psi_{\infty}}^{\pm 1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C.$
    • $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi_{\infty}^{\pm 1}|^{2}_{T}T^{n}\leq C$.
@@ 3. $(3)$
3.3 Existence of $T$-adapted HYM metrics
Notation 3.15.
We use the notations in Notation 3.2. Let $\Psi_{\infty}^{\pm 1}$ be a smooth $h_{0}$-hermitian endomorphism of $E|_{X\setminus D}$ constructed in Theorem 3.14.
@@ 3.3 Existence of $T$-adapted HYM metrics
We use notations in Notation 3.15. Then the following holds.
@@ 1. $(1)$
    The smooth hermitian metric $h_{\infty}:=h_{0}\Psi_{\infty}^{2}$ on $E|_{X\setminus D}$ defines a $T$-adapted HYM metric on $E$ which satisfies
    • $\sup_{X\setminus D}|s_{D}\Psi_{\infty}^{\pm 1}|_{h_{0}\otimes h_{0}\otimes h_{D}}\leq C$ and
    • $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi_{\infty}^{\pm 1}|^{2}_{h_{0}\otimes h_{0}\otimes h_{D}^{m+1},T}T^{n}\leq C$.
@@ 2. $(2)$
    Furthermore, the $T$-HYM constant $\lambda$ of $h_{\infty}$ satisfies the following:
    $\lambda=\frac{1}{\alpha^{n}{\rm{rk}}E}\int_{X\setminus D}\sqrt{-1}{\rm{Tr}}(F_{h_{\infty}})\wedge T^{n-1}=\frac{1}{\alpha^{n}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
@@
Since $h_{i}=h_{0}\Psi_{i}^{2}$ is an $\omega_{i}$-HYM metric on $E$, it locally satisfies
$\sqrt{-1}\Lambda_{\omega_{i}}\bar{\partial}(\partial h_{i}h_{i}^{-1})=\lambda_{i}{\rm{id}}$ (3.34)
where $\lambda_{i}$ is the $\omega_{i}$-HYM constant of $h_{i}$ described as
$\lambda_{i}=\frac{1}{\{\omega_{i}\}^{n}}\frac{c_{1}(E)\cdot\{\omega_{i}\}^{n-1}}{{\rm{rk}}E}.$ (3.35)
Since $\Psi_{i}^{\pm 1}$ converges to $\Psi_{\infty}^{\pm 1}$ in $C^{1,\alpha}$-topology on each relatively compact open subset in $X\setminus D$, the equation (3.34) also weakly converges in $X\setminus D$. Hence, on each small compact subset in $X\setminus D$, we have that $h_{\infty}$ satisfies
$\sqrt{-1}\Lambda_{T}F_{h_{\infty}}=\sqrt{-1}\Lambda_{T}\bar{\partial}(\partial h_{\infty}h_{\infty}^{-1})=\lambda{\rm{id}}$ (3.36)
in the sense of currents. Since (3.36) is the $T$-HYM equation, the constant $\lambda$ is the $T$-HYM constant of $h_{\infty}$. By (3.35) and the integral of (3.36) over $X\setminus D$ against $T^{n}$, we obtain
$\lambda=\frac{1}{\alpha^{n}{\rm{rk}}E}\int_{X\setminus D}\sqrt{-1}{\rm{Tr}}(F_{h_{\infty}})\wedge T^{n-1}=\frac{1}{\alpha^{n-1}}\frac{c_{1}(E)\cdot\alpha^{n-1}}{{\rm{rk}}E}.$
The $T$-adapted condition of $T$ is a direct consequence of Theorem 3.14. ∎
By Theorem 3.14, we obtain the following.
@@
    $\int_{X\setminus D}|s_{D}^{m+1}\partial^{\nabla_{h_{0}}}\Psi^{\pm 1}|^{2}T^{n}\leq C.$
  3. $(3)$
    The $T$-HYM constant of $h$ satisfies